OpenStudy (anonymous):

solve the equation. x^2 + 4x - 60 = 0

OpenStudy (anonymous):

Let f(x) = x^2 - 4x - 60. The graph of f(x) = 0 is a parabola which opens up and has x-intercepts x = -6 and x = 10. The points satisfying f(x) <= 0 are the points which are between the intercepts, so -6 <= x <= 10.

OpenStudy (anonymous):

factor and you get; (x-6)(x+10)=0, so either x-6=0 or x+10=0 so solving for x you get; x=6,x=-10

OpenStudy (anonymous):

thank you sooooo much both of you

OpenStudy (anonymous):

give us medals?;)

OpenStudy (anonymous):

how ido i give you medals im trying to find out how to do it this is my first time

OpenStudy (anonymous):

click next name good answer

OpenStudy (anonymous):

okay thank you i did it im 17 years old and i struggle alot in math so you really helped me alot on that problem i was stuck on for like 20 minutes

OpenStudy (anonymous):

i attend Detroit School of the Fine And Performing Arts were in my school is excellent but i just cant get the math its too hard

OpenStudy (anonymous):

IIM 17 and do online school i also blow at math